у х х хThe volume of the triangular block is 4 cubic inches.

What is the approximate length of y? Round to the nearest tenth of an inch.

1.4 in.

2.0 in.

2.8 in.

3.5 in.
Answer #1

Volume of box is (x*x*x/2)=4

= x^3=8

X=2.

Now the relationship between x and y , by pythagoras

X^2+x^2=y^2

Y^2= 2*(x)^2

Y^2= 2×2^2

Y^2= 8

Y=2√2

Y approximately equal to 3 inch.

Hopefully this will be helpful for you thanks please upvote it